Complete each ordered pair so that it satisfies the given equation:
1. f(x)=x^2 (4, ) ( ,9)

Answer by ankor@dixie-net.com(22740)

Complete each ordered pair so that it satisfies the given equation:
1. f(x) = x^2 (4, ) ( ,9)
y = f(x)
:
y = x^2
x = 4
therefore
y = 4^2
y = 16
1st ordered pair: 4,16
:
y = 9
x^2 = 9
therefore
x = 3
2nd ordered pair: 3, 9